# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

### v3

@@ 3.2. Defining Sources of Loss
The four steps above identify four sources of loss representing possible reasons why observed human performance in an experiment might not match the normative standard.
Prior loss (1). The agent must obtain prior beliefs that match the prior beliefs the researchers assume in analyzing the problem. Prior loss is the loss in performance due to the difference between the agent’s prior beliefs and those used by the researchers to calculate the normative standard. Explicit specification of priors is rare in decision studies used to evaluate interfaces in practice. The consequences of this vary depending on the nature of the signaling strategy. When the signal is fully revealing of the state (i.e., the posterior is a point mass indicating the state with certainty), or directly provides the posterior probability of the state, there is no need for the prior. However, whenever the signal cannot directly provide the probability of the state, the prior must be communicated for a well-defined decision problem.
@@
An important implication of the proper scoring rule distinction for belief reports is that in general we should not expect an agent to truthfully report their beliefs when it is not in their best interest to do so. Consequently, when non-proper rules are used for belief tasks, we cannot interpret a study participant who does not report the appropriate beliefs after examining the signal as demonstrating bias. Such “deviation” from what the researcher might expect (honest reporting) is not technically a loss, it is the expected behavior under that rule. In practice, researchers often use non-proper scoring rules when eliciting predictions; see, e.g., (Gneiting, 2011) for examples from finance where commonly used rules allow forecasters to score better by not reporting the true conditional mean of the target random variable.
@@
Misunderstanding the decision problem. A potential contributor to all four sources of loss above is the agent misunderstanding the decision task, including the scoring rule, the statespace, and/or the action space. Even when all necessary components of a decision problem are carefully communicated, the participant might not read the instructions carefully, or might not understand them. Consequently, we should think of any tests we run to validate hypotheses about biases stemming from particular sources of performance loss as tests of the joint hypothesis comprised of the target hypothesis and a hypothesis that participants understood the task.
3.3. Examples
Researchers are limited in their ability to interpret responses to decision experiments as biased when distinct sources of performance loss that account for deviation from the normative standard cannot be conceived. We use several examples inspired by the human-centered AI and visualization literature to demonstrate why drawing conclusions about biased behavior is misleading in such scenarios.
3.3.1. AI-Assisted Recidivism Decisions
@@
We first elaborate the recidivism decision-making study mentioned earlier. Imagine the study is designed to investigate whether people over-rely on AI assistance when deciding whether to release (i.e., grant parole to) incarcerated defendants. The researchers use a standard train-validate-split approach on a large labeled historical data set to train a model that predicts whether or not a person will recidivate. Each defendant in the dataset is represented by a set of feature values, , which comprise part of the signal shown on each trial. Study participants are randomly assigned to either a control condition where they see the feature values along with only the model’s prediction and what is described as “the model’s confidence in its prediction, ranging between 50% and 100%,” or an AI explanation condition where they are also given additional information about how features in the data contribute to the model’s predictions, e.g., generated by LIME (Ribeiro et al., 2016). The model confidence score that is provided is the probability of a correct prediction conditional on the feature values, .
@@
Participants are told the AI model was trained on prior defendants whose status is known. They are told that the model has an accuracy of (representing the combined false positives and false negatives) on a sample of instances drawn from the test set which comprise the trials in the study.
@@
Each participant makes parole decisions for 20 unique defendants. On each trial they are asked to choose whether or not to release the defendant (). Rather than the scoring rule given in Equation 5, imagine a simpler version that assigns different types of errors equivalent costs. Participants are given a base reward of , but they are told they will lose from this initial endowment for every trial where they either release a person who will re-offend or continue to incarcerate a person who will not re-offend. Participants are not told how often to expect recidivism; i.e., the prior is not endowed. Participants are not given interim feedback on their decisions.
@@
After collecting data from 500 participants, the researchers model the percentage of trials where the participants incorrectly follow the AI’s prediction. They find a statistically significant difference between the AI explanation condition and the control condition. The researchers interpret this as evidence that AI explanations promote over-reliance.
@@
Interpretation of Results. To assess what can be said about performance loss, we first consider whether the problem is well-defined: could a rational agent could use the provided information to select the utility-maximizing decision?
@@
The action space ( for grant parole, for deny parole) is conveyed to participants, as well as the statespace ( for the defendant re-offending, for not). They also have knowledge of the scoring rule that assigns a score (reward amount) given the realized state and their decision.
@@
How would a rational participant form beliefs about the posterior probability of the state using Equation 3? Participants are not provided with the prior probability of the state. This would be acceptable if the signal provided them with posterior beliefs about the uncertain state. However, the signal does not provide the required posterior probability: recall that participants receive the probability that the model is correct conditional on the features, , but not conditional on the model’s prediction. They therefore cannot obtain sufficient information from the data-generating model to guide their choice of action.
@@
However, the careful reader may also have noticed another barrier to defining bias for the problem: the model’s conditional confidence is described as simply “the model’s confidence in its prediction”, making it ambiguous from the participant’s perspective whether it is an unconditional or conditional. Given that the normative response is not well-defined in the unconditional case even with the prior, this description would need specify that the confidence is conditional on the predicted label to eliminate ambiguity.
In the case where the study presented only , not e.g., when the model predicts 1, what additional information would equip the agent to optimize? If instead of the overall model accuracy over the stimuli, the experimenters provided information about the model’s error as separated rates of false positives and false negatives, the participant would be equipped with enough information to infer in Equation 3.
It is worth noting that decision problems do not necessarily have to depend on small, discrete action spaces. Imagine another amendment to the recidivism study where instead of asking participants for a parole decision, the researchers ask them to report the probability of recidivism (see, e.g., (Fogliato et al., 2021; Green and Chen, 2019)). Assuming the above changes are made to the design to equip participants with information to form posterior beliefs, participants could in principle report the optimal beliefs. However, if the scoring rule for a decision problem where the action space is over beliefs is not proper, we should not expect the agent to respond with their true beliefs. Optimization loss might be low, but this is not useful to the experimenters’ research questions if agents respond with a distortion of their true beliefs.
3.3.2. Election Forecast Displays
@@
Participants are told they will win a given base reward amount plus an additional “for each trial where their candidate wins a simulated version of the election that uses the forecast model to predict how other people will vote.” The researchers generate outcomes by randomly drawing a bivariate vote share outcome from the predicted distribution under the model and generating the uncertain state (taking the value 0 or 1 for lose or win) by assuming the higher vote share candidate wins. Participants are not given feedback after each trial.
After collecting results for several hundred participants, the researchers fit a linear model that estimates the sensitivity of the vote decision to the closeness of the race, defined as the difference between the predicted vote share of the first candidate and 0.5. The researchers postulate that a better display induces more sensitivity. Using statistical significance testing, they reject the null hypothesis of no difference between the estimated slopes for the interaction between display type and race closeness across display types, using posthoc comparisons to declare that the animated graphic leads to better voting decisions because it induces the most sensitivity.
Interpretation of Results. To understand why the researchers’ conclusions might be misleading as a prediction of real world use of election displays, consider the decision problem from the perspective of the participants. The action space ( for vote, for not vote) is conveyed to participants, and the task description also gives participants the statespace ( for their candidate winning, for not). Participants can infer the probability of their candidate winning from the signal showing the model predicted vote share.
What do participants know about the realized outcome of the election that the researchers describe simulating? The researchers described the use of the forecast model to simulate voting behavior, and in implemention, do not account for the participant’s vote in generating the outcome. Consequently, whether a participant votes is irrelevant to the payoff-relevant state under the scoring rule. However, the description provided to participants is ambiguous about whether the participant’s vote counts: some participants might assume, based on their knowledge of real-world elections, that it does. Hence performing well in the experiment requires being able to guess what the researchers have in mind, rather than behaving rationally in the context of the experiment.
@@
However, the decision problem does not admit a well-defined normative response until we address the ambiguity in the description of the data-generating model. Only once participants are given a transparent description of how the realized outcome is generated from the model can we conceive of prior loss as well as updating loss, both of which should be zero if the problem is understood because the prior does not affect the judgment. We can conceive of receiver loss as the loss from participants not being able to extract the information from the signal, and optimization loss as the loss from participants not being able to choose the optimal action under their beliefs.
Of course, if we amend the description to inform participants of how the election outcome is actually generated, we should expect participants who understand the problem to never vote. Rather than informed political engagement, the proportion of participants who vote (e.g., in each treatment condition) would be better described as reflecting the level of confusion about the task. Even if the researchers did not change the description of how the realized state is generated, but designed the study to give participants feedback on the realized state after each decision, then the participants might in theory learn that their vote does not matter, which works against the researchers’ goals of presenting a situation where people could feel compelled to vote.
@@
3.4. Goals & Incentives in the Lab vs World
Imagine that the researchers who designed the voting task intentionally did not include the vote in the scoring rule, and were intentionally ambiguous about the role of the vote in the data-generating model. Because the utility of voting in a real-world election is typically left implicit and likely to vary between people, the researchers might think that their results would therefore be more representative of real world voting behaviors.
Consequently, when people are in an experiment, we should not expect the analogous real world goals and beliefs for that kind of task to apply. For example, people might take actions in the real world for intrinsic value—e.g., voting in order to feel like a good citizen, or paying close attention to a defendant’s profile because the decision of whether to grant parole or not is expected to have a tremendous impact on that person’s life. It is difficult to motivate people to take actions based on intrinsic value in an experiment, unless the experiment is designed to elicit and describe those intrinsic values, in which case the normative framework here no longer applies. Hence, the idea that as experimenters we can skip fully describing the decision problem to our participants while still retaining the ability to identify their behavior as faulty or biased arises from a failure to recognize our fundamental uncertainty about how the experimental context relates to the real world.
4. Applying the Framework to HCAI Studies
@@
We use a sample of recent evaluative studies of AI-assisted decisions to estimate how often researchers draw conclusions about flaws in human decisions in the absence of a well-defined decision problem. We sampled 46 of the studies surveyed by Lai et al. (2023) in their overview of empirical human-subjects studies on human-AI decision-making. According to their inclusion criteria, the sample contains only evaluative human-subject studies targeting a decision task in the context of classification or regression published in ACM or ACL conference venues between 2018 and 2021.
@@
Codes. In coding the studies, we focused on three aspects of the study design and interpretation. First, we identified whether our framework is applicable: Does the task given to participants admit a ground truth state and induce uncertainty about the state, at least across participants. Next, we identified whether the authors draw conclusions about flaws or performance loss in human decisions from their results, i.e., is the study evaluating human decision-making in some way. We identified whether the study targeted a well-defined problem; i.e., whether participants in the study were provided with sufficient information for a rational agent to best respond to the task. Finally, we noted whether the researchers reported conclusions using a different scoring rule than the one provided to participants, and whether participants were incentivized using a flat scoring rule (i.e., a rule that does not distinguish between responses based on their distance to the ground truth). Note that studies that pay participants a flat reward require that researchers report conclusions using a different scoring rule. Our coding guide is in Appendix 7.2.
@@
In some cases, coding was challenging because the paper did not provide much detail on what exactly was communicated to participants, nor provide the study interface as a supplement. We note such cases in our coding table (Appendix 7.3).
@@
Two authors participated in the coding. After formulating the coding scheme, the authors divided the sampled studies roughly evenly, each coding independently, then discussing all challenging examples to ensure consensus. We provide the list of sampled studies with codes as Appendix.
@@
We observed that even when authors gave participants sufficient information to identify the best response, they rarely compared the observed performance to the best attainable performance. For example, multiple papers defined a scoring rule to incentivize participants to make accurate decisions, but evaluated the responses using a simpler accuracy scale, implying that 100% accuracy was obtainable even when it would not be to a perfectly-perceiving rational agent,and potentially misinterpreting the performance of participants guided by a different scoring rule. In some cases, participants were also given sufficient information to optimize, but it did not accurately reflect the true data-generating model. The most common scenario was when the researchers gave participants information about the model’s held-out accuracy but the participants experienced instances where the model had a different accuracy. When the study design includes feedback, this leads to a contradiction between the instructions and what participants observe which makes it difficult to draw evaluative claims about behavior.

### v5

@@ 2.4. Defining Sources of Loss
The four steps above identify four sources of loss representing possible reasons why observed human performance in an experiment might not match the normative standard. A final source of loss that must be considered when dealing with human (as opposed to hypothetical rational) agents is misunderstanding of the decision problem.
Prior loss (1). The agent must obtain prior beliefs that match the prior beliefs the researchers assume in analyzing the problem. Prior loss is the loss in performance due to the difference between the agent’s prior beliefs and those used by the researchers to calculate the normative standard.
@@
Misunderstanding the decision problem. A potential contributor to all four sources of loss above is the agent misunderstanding the decision task, including the scoring rule, the statespace, and/or the action space. Even when all necessary components of a decision problem are carefully communicated, the participant might not read the instructions carefully, or might not understand them. Consequently, any tests that the researcher runs to validate hypotheses about biases stemming from particular sources of performance loss should be thought of as tests of the joint hypothesis comprised of the target hypothesis and the hypothesis that participants understood the task.
3. Applying the Normative Decision Framework in Practice
We outline the scope of the above definitions in terms of what kinds of human subjects studies they apply to, and describe the primary implications for researchers seeking to evaluate human decisions. We discuss considerations for the design of the signal and scoring rule.
3.1. Scope
Our definition of a decision problem and corresponding standard for optimal behavior is intended for controlled evaluation of human behavior, a.k.a. normative decision research. Behavioral data (which may be produced by humans or by simulation) is collected under controlled conditions, with the goal of learning about behavior induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank different assistive elements by human performance (e.g, different visualizations (Correll et al., 2018) or AI explanation strategies (Liu et al., 2021)), or to test a hypothesis about how humans make decisions or what will help them do better (e.g., cognitive forcing functions will improve AI-assisted decisions (Buçinca et al., 2021)). Critically, such evaluative studies require that ground truth can be identified for any hypothetical states of the world against which study participants’ decisions are evaluated.
It is important to recognize that even if the researchers do not think about the elements of their experiment in terms of the components identified above (i.e., statespace, data-generating model, signaling policy, action space, scoring rule), the design of their study still induces a decision problem whenever two simple conditions hold (Wu et al., 2023). First, there must be some uncertainty about the state (i.e., there is more than one data scenario used to generate stimuli)111A single trial study where all participants view the same signal would not qualify., and second, the researchers believe it is possible to provide a scoring rule that is consistent across participants. This requirement eliminates, for example, studies of emotional responses.
3.2. Implications for communicating decision tasks to study participants
The primary motivations for using the above framework are epistemological in nature: they concern what knowledge can be gained from the results of a study. Our primary claim in this work is that to evaluate human decision-making, an experiment must provide the participant with enough information to in principle identify the normative decision, which is used to judge their behavior. In other words, does the experiment give the participant enough information to align their understanding of the decision problem with its normative interpretation? Researchers must communicate to participants the action space, state space, scoring rule, and sufficient information about the data-generating model for participants to in theory arrive at the posterior beliefs against which they are evaluated.
When the experiment provides the participant sufficient information, we can characterize forms of bias as loss in decision performance according to the four possible sources described in Section 2.4. When an experiment does not provide sufficient information for the participant to solve the task in principle, there is the potential for multiplicity concerning the data-generating model. Such multiplicity has implications for interpreting the study results, as we demonstrate through examples in Section 2.3.
3.3. Designing signals and feedback
There is often more than one way for a study to provide sufficient information for a rational decision-maker to identify the normative decision. Experiment designs can vary in the amount of information they provide participants about or how they provide this information, provided enough information is conveyed for a rational agent to identify the best response.
For example, for many experiments, it is critical to convey or an approximation of as the prior in order for participants to be able to update these beliefs upon viewing signals. However, for some experiments, it is instead possible for the signal to directly inform of the posterior probability of the state. For example, in a visualization study that asks users to decide whether or not they want to purchase a piece of special equipment expected to improve their score in a game (e.g., (Hofman et al., 2020; Kale et al., 2020)), a visualization of the player’s expected score distribution with and without the equipment (from which the realized outcome is drawn depending on their choice) conveys the posterior probability of the state directly.
Some experiment designs support learning about by providing participants feedback informing them of the realized state after each decision trial. When feedback is provided, we would expect the Bayesian rational decision-maker to update their beliefs according to the information provided on the joint distribution between signals and states. Imagine that participants in the flight booking study see feedback (i.e., are told the realized state, whether or not the flight was delayed) after each decision they make. Assuming the signal is not revealing of the posterior probability of the state (delayed or not), we would naturally expect the participants to learn about from this feedback as they complete trials, affecting their posterior beliefs (Equations 1 and 2). A challenge in using this approach without providing other information about (i.e., through signals and the prior) is that how much a participant can learn from any single trial is quite limited. Hence, such designs tend to require many trials.
3.4. Choice of Scoring Rule
Whether the researcher uses the same scoring rule for both incentives and evaluation versus two different rules has important implications for interpreting the results. For the flight booking example, if the researchers analyzed the collected responses using the above rule (i.e., reporting performance in payoff space) in addition to communicating it to participants for incentivization, then the utility associated with performance is defined by a single scoring rule, and the interpretation of a “good” decision is consistent across incentivization and evaluation, which is clearly beneficial for interpreting the results.
One way around this problem when using two different rules is to make sure that the rule used for incentivization has certain properties. Specifically, among scoring rules for beliefs, we can distinguish those that are proper scoring rules: scoring rules with for which the optimal action is to predict the true distribution, i.e., . When a proper scoring rule is used, the researcher can assume that an agent who understands the rule properly will report their true beliefs, as it is not possible to score higher by reporting some alternative beliefs. These beliefs can then be plugged into a different scoring rule without the risk of misinterpretation.
An important implication of the proper scoring rule distinction for belief reports is that in general we should not expect an agent to truthfully report their beliefs when it is not in their best interest to do so. Consequently, when non-proper rules are used for belief tasks, we cannot interpret a study participant who does not report the appropriate beliefs after examining the signal as demonstrating bias. Such “deviation” from what the researcher might expect (honest reporting) is not technically a loss, it should be the expected behavior under that rule.
Applying the construction in Equation 7 to a binary decision gives a threshold rule where for each state, all reported probabilities below a threshold receive one score and those above receive another score. Since the payoff depends only weakly on the belief, such scoring rules may be less sensitive to the agents’ understanding of the signal. For this reason, including a decision problem that directly also elicits beliefs using a strictly proper-scoring rule is a good idea whenever eliciting responses using an action space that is not over beliefs, as it enables plugging the elicited beliefs into other scoring rules.222See, e.g., Wu et al. (2023)’s analysis of (Kale et al., 2020).
The choice of possible scoring rules depends on the nature of the action space. In general, a coarser action space (i.e., with fewer choices available to participants) will produce less informative responses for the purpose of resolving observed performance relative to a normative standard. If, for example, the researcher designed the flight booking study to also elicit the agent’s prediction of the probability of delay, each trial would provide more information, and they could employ a proper scoring rule such as squared loss (i.e., Brier or quadratic rule) or log score to score the accuracy of the probabilistic beliefs. Different proper scoring rules imply different expectations about the set of possible decisions (Levinstein, 2017). See Li et al. (2022) for a discussion of scoring rules that have a high information value for the rational agent.
4. Examples
We use several examples inspired by the human-centered AI and visualization literature to demonstrate why drawing conclusions about biases in human behavior is misleading in scenarios where sufficient information for a rational agent to optimize is not provided. We discuss how the example study designs could be improved to align with the definitions in Section 2.
4.1. AI-Assisted Flight Booking
@@
We first elaborate the flight booking study mentioned earlier. Imagine the study is designed to investigate whether people over-rely on AI assistance when deciding whether to book a flight or not. The researchers use a standard train-validate-split approach on a large labeled historical data set to train a model that predicts whether or not a flight will be delayed. Each flight in the dataset is represented by a set of feature values, , which comprise part of the signal shown on each trial. Study participants are randomly assigned to either a control condition where they see the feature values along with only the model’s prediction and what is described as “the model’s confidence in its prediction, ranging between 50% and 100%,” or an AI explanation condition where they are also given additional information about how features in the data contribute to the model’s predictions, e.g., generated by LIME (Ribeiro et al., 2016). The model confidence score that is provided is an approximately calibrated probability333A confidence function is -calibrated with respect to a sample if there exists some , with , such that for any : (Corvelo Benz and Rodriguez, 2024). In other words, the probabilities are approximately calibrated with respect to a sample when for some subset at least times the size of , the difference between the stated and real probability is less than or equal to . of a correct prediction conditional on the feature values, .
@@
Participants are told the AI model was trained on prior flights whose delay status is known. They are told that the model has an accuracy of (representing the combined false positives and false negatives) on a sample of instances drawn from the test set which comprise the trials in the study.
@@
Each participant makes booking decisions for 20 unique flights. On each trial they are asked to choose whether or not to book the flight (). Rather than the scoring rule given in Equation 6, imagine a simpler version that assigns different types of errors equivalent costs. Participants are given a base reward of , but they are told they will lose from this initial endowment for every trial where they either book a flight that will be delayed or do no book a flight that will not be delayed. Participants are not told how often to expect delays; i.e., the prior is not endowed, nor is . Participants are not given interim feedback on their decisions.
@@
After collecting data from 500 participants, the researchers model the percentage of trials where the participants incorrectly follow the AI’s advice. They find a statistically significant difference between the AI explanation condition and the control condition. The researchers interpret this as evidence that AI explanations promote over-reliance.
@@
Interpretation of Results. To assess what can be said about performance loss, we first consider whether the problem is well-defined: could a rational agent use the provided information to select the utility-maximizing decision?
@@
The action space ( for book, for not book) is conveyed to participants, as well as the statespace ( for the flight being delayed, for not). Through the information they are given about the task payouts, they also have knowledge of the scoring rule that assigns a score (reward amount) given the realized state and their decision.
@@
How would a rational participant form beliefs about the posterior probability of the state using Equation 1? Participants are not provided with information about the prior probability of the state. This would be acceptable if the signal provided them with posterior beliefs about the uncertain state. However, the signal does not provide the required posterior probability: recall that participants receive the estimated probability that the model is correct conditional on the features, , but not conditional on the model’s prediction. They therefore cannot obtain sufficient information from the data-generating model to guide their choice of action.
@@
However, the careful reader may also have noticed another barrier to defining bias for the problem: the model’s conditional confidence is described as simply “the model’s confidence in its prediction”, making it ambiguous from the participant’s perspective whether it is conditional on the features and model prediction or only one of the two. The normative response is not well-defined if the probability is not conditional on both, unless the researchers also communicate or an approximation to participants before they begin the decision trials. Hence the model probability description would need to be amended to specify that the confidence is conditional on the predicted label to eliminate ambiguity.
It is worth noting that decision problems do not necessarily have to depend on small, discrete action spaces. Imagine another amendment to the flight booking study where instead of asking participants for a booking decision, the researchers ask them to report the probability that the flight is delayed . Assuming the above changes are made to the design to equip participants with information to form posterior beliefs, participants could in principle report the optimal beliefs. However, if the scoring rule for a decision problem where the action space is over beliefs is not proper, we should not expect the agent to respond with their true beliefs. Optimization loss might be low, but this is not useful to the experimenters’ research questions if agents respond with a distortion of their true beliefs.
4.2. Election Forecast Displays
@@
Participants are told they will win a given base reward amount plus an additional “for each trial where their candidate wins a simulated version of the election that uses the forecast model to predict how people will vote.” The researchers generate outcomes by randomly drawing a bivariate vote share outcome from the predicted distribution under the model and generating the uncertain state (taking the value 0 or 1 for lose or win) by assuming the higher vote share candidate wins. Participants are not given feedback after each trial.
After collecting results for several hundred participants, the researchers fit a linear model that estimates the sensitivity of the vote decision to the closeness of the race, defined as the difference between the predicted vote share of the first candidate and 0.5. The researchers postulate that a better display induces more sensitivity. Using statistical significance testing, they reject the null hypothesis of no difference between the estimated slopes for the interaction between display type and race closeness across display types. Using posthoc comparisons, they declare that the animated graphic leads to better voting decisions because it induces the most sensitivity.
Interpretation of Results. To understand why the researchers’ conclusions might be misleading as a prediction of real world use of election displays, consider the decision problem from the perspective of the participants. The action space ( for vote, for not vote) is conveyed to participants, and the task description also gives participants the statespace ( for their candidate winning, for not). Participants can infer the probability of their candidate winning from the signal showing the model predicted vote share. Because the signal provides the posterior probability of the state, no prior information needs to be communicated.
What do participants know about the realized outcome of the election that the researchers describe simulating? The researchers described the use of the forecast model to simulate voting behavior, and in implemention, do not account for the participant’s vote in generating the outcome. Consequently, whether a participant votes is irrelevant to the payoff-relevant state under the scoring rule. However, the description provided to participants is ambiguous about whether the participant’s vote counts. Some participants might assume, based on their knowledge of real-world elections, that it does. Hence performing well in the experiment requires being able to guess what the researchers have in mind, rather than behaving rationally in the context of the experiment.
@@
However, the decision problem does not admit a well-defined normative response until we address the ambiguity in the description of the data-generating model. Only if participants are given a transparent description of how the realized outcome is generated from the model can we conceive of prior loss as well as updating loss, both of which should be zero given the signal design. We can conceive of receiver loss as the loss from participants not being able to extract the information from the signal, and optimization loss as the loss from participants not being able to choose the optimal action under their beliefs.
Of course, if we amend the description to inform participants of how the election outcome is actually generated, we should expect participants who understand the problem to never vote. Rather than informed political engagement, the proportion of participants who vote (e.g., in each treatment condition) could be described as reflecting the level of confusion the participants had about the task. Or, if the researchers did not change the description of how the realized state is generated, but designed the study to give participants feedback on the realized state after each decision, then the participants might in theory learn that their vote does not matter. This would work against the researchers’ goals of presenting a situation where people could feel compelled to vote.
@@
5. Applying the Normative Framework to HCAI Studies
To illustrate the applicability of our definition, we present a meta-study in which we analyzed a sample of recent evaluative studies of AI-assisted decisions to estimate how often researchers draw conclusions about flaws in human decisions without a well-defined decision problem.
@@ 5. Applying the Normative Framework to HCAI Studies
Sample. We sampled 46 of the studies surveyed by Lai et al. (2023) in their overview of empirical human-subjects studies on human-AI decision-making. According to their inclusion criteria, the sample contains only evaluative human-subject studies targeting a decision task in the context of classification or regression published in ACM or ACL conference venues between 2018 and 2021.
@@
Codes. In coding the studies, we focused on three aspects of the study design and interpretation. First, we identified whether our framework is applicable: Does the task given to participants admit a ground truth state and induce uncertainty about the state, at least across participants? Next, we identified whether the authors draw conclusions about flaws or performance loss in human decisions from their results, i.e., is the study evaluating human decision-making in some way? We identified whether the study targeted a well-defined problem; i.e., whether participants in the study were provided with sufficient information for a rational agent to best respond to the task. Finally, we noted whether the researchers reported conclusions using a different scoring rule than the one provided to participants, and whether participants were incentivized using a flat scoring rule (i.e., a rule that does not distinguish between responses based on their distance to the ground truth). Note that studies that pay participants a flat reward require that researchers report conclusions using a different scoring rule.
@@
In some cases, coding was challenging because the paper did not provide much detail on what exactly was communicated to participants, nor provide the study interface as a supplement. We noted such cases.
@@
Two authors participated in the coding. After formulating the coding scheme, the authors divided the sampled studies roughly evenly, each coding independently, then discussing all challenging examples to ensure consensus. The list of coded studies, our coding guide, and full coding results are available as an Appendix.
@@
We observed that even when authors gave participants sufficient information to identify the best response, they rarely compared the observed performance to the best attainable performance. For example, multiple papers defined a scoring rule to incentivize participants to make accurate decisions, but evaluated the responses using a simpler accuracy scale, implying that 100% accuracy was obtainable even when it would not be to a perfectly-perceiving rational agent. As described above, this misinterprets the performance of participants guided by a different scoring rule. In some cases, participants were also given sufficient information to optimize, but it did not accurately reflect the true data-generating model. The most common scenario was when the researchers gave participants information about the model’s held-out accuracy but the participants experienced instances where the model had a different accuracy. When the study design includes feedback, this leads to a contradiction between the instructions and what participants observe which makes it difficult to draw evaluative claims about behavior.